Dul Date: Wednesday. August 1, 2018 Student Name: 11. The system is released from rest. The 4-kg mass slides on the smooth horizontal surface. ing the principle of work and energy, determine the magnitude of the velocity of the By us masses when the 20 kg mass has fallen 1 m. 4 kg 20 kg
Answer #1

here,

m1 = 4 kg

m2 = 20 kg

h = 1 m

let the final velocities be v

using conservation of energy

change in potential energy = kinetic energy gained

m2 * g * h = 0.5 * ( m1 + m2) * v^2

20 * 9.81 * 1 = 0.5 * ( 20 + 4) * v^2

solving for v

v = 4.04 m/s

the magnitude of final velocities is 4.04 m/s